the perimeter of a right triangle is 90 feet. the ratio of the legs is 6:8. what is the length of the longest leg of the triangle
Answer by checkley77(12844)

Seeing as you have a right triangle with 2 sides=6 & 8 then:
6^2+8^2=c^2
36+64=c^2
100=c^2
c^2=100
c=sqrt100
c=10
Therefore:
6n+8n+10n=90 [n=the ratio multiplier]
24n=90
n=90/24
n=3.75 this is the multiplier for the ratio values.
6*3.75=22.5
8*3.75=30
10*3.75=37.5
Proof:
22.5+30+37.5=90
90=90